OpenStudy (anonymous):

HIJK IS A RECTANGLE. FIND THE VALUE OF X AND THE LENGTH OF EACH DIAGONAL hj=19+2x and ik=3x+22

OpenStudy (anonymous):

Diagonals of a rectangle are of equal length

OpenStudy (anonymous):

how does hj=to x

OpenStudy (anonymous):

oh ok that's fine

OpenStudy (anonymous):

I still don't understand

hartnn (hartnn):

@bubblez20133 hi :) \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) as @Xavier said, 'Diagonals of a rectangle are of equal length' means we equate hj=ik so, 19+2x=3x+22 can you find 'x' from here ?

OpenStudy (anonymous):

I can try and thank you I appreciate you welcome.

OpenStudy (anonymous):

i got x=3

hartnn (hartnn):

very close, its actually, x=-3

OpenStudy (anonymous):

so why is it negative 3

hartnn (hartnn):

19+2x=3x+22 subtracting 2x from both sides, 19 = x +22 subtracting 22 from both sides, x = 19-22 = -3

OpenStudy (anonymous):

ohhhh thanks you are so helpful

hartnn (hartnn):

welcome ^_^